what is the slope of 4x=7y=16
Found 2 solutions by stanbon, rfadrogane:
Answer by stanbon(75887)

what is the slope of 4x=7y=16
----
One to many equal signs.
Cheers,
Stan H.

Answer by rfadrogane(214)

what is the slope of 4x-7y=16
Sol'n:
4x-7y=16
7y=4x-16, divide 7 on both side
y=(4/7)x-16/7
so, from y = mx+b, m is the slope
thus, from the equation above.
m = 4/7 ----answer